# Multichannel Conflict-Avoiding Codes for Expanded Scenarios

### v1

@@ IV-B Codeword Length $L=p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{n}^{r_{n}}$
This subsection is devoted to generalize the following result, which is due to [5, Theorem 10].
@@
Let $w$ be a positive integer, and let $p_{1}<p_{2}<\cdots<p_{n}$ be primes satisfying $p_{1}\geq 2w-1$. If there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $m_{i}$ codewords for each $i$, then for any positive integers $r_{1},\ldots,r_{n}$, there exists a code $\mathcal{C}\in\mathrm{CAC^{e}}(L,w)$ with
@@
$g\in\mathcal{S}_{r_{i}}(\Gamma_{i}),x_{i}\in\mathbb{Z}_{p_{i}}^{r_{i}}\text{ for }i<j\leq n\}.$
@@
Suppose to the contrary that $d^{*}(S_{a})\cap d^{*}(S_{b})\neq\emptyset$, where $a=(0,\ldots,0,g,x_{i+1},\ldots,x_{n})$ and $b=(0,\ldots,0,h,y_{i+1},\ldots,y_{n})$ for some $g,h\in\mathcal{S}_{r_{i}}(\Gamma_{i})$. Assume $j(0,\ldots,0,g,x_{i+1},\ldots,x_{n})=k(0,\ldots,0,h,y_{i+1},\ldots,y_{n})$ for some $j,k\in\{\pm 1,\pm 2,\ldots,\pm(w_{i}-1)\}$. If $j=k$, then $g=h$ and $x_{t}=y_{t}$ for $i+1\leq t\leq n$, namely, $a=b$. So, we consider $j\neq k$ in what follows. Let us focus on the $i$-th component, namely, $jg=kh$. Since $p_{i}\geq 2w-1$, by Proposition 3, one has $jg\in\mathsf{L}_{t}$ for every $j\in\{\pm 1,\pm 2,\ldots,\pm(w-1)\}$ if $g\in\mathsf{L}_{t}$. Here we use $\mathsf{L}_{t}$ to denote the $t$-th layer in the $p_{i}$-ary representation of elements of $\mathbb{Z}_{p_{i}^{r_{i}}}$ for notatinoal convenience. It follows that $jg\neq kh$ for any $j,k\in\{\pm 1,\pm 2,\ldots,\pm(w-1)\}$ whenever $g$ and $h$ are in distinct layers. Therefore, we only need to consider the case when both $g$ and $h$ are in the same layer, say $\mathsf{L}_{t}$ for some $t$. Observe that $j,k\in\mathsf{L}_{0}$ due to $p_{i}\geq 2w-1$. By Proposition 3 again, $j\cdot g_{t}=k\cdot h_{t}$ (mod $p_{i}$). If $g_{t}\neq h_{t}$, by the assumption that $g_{t},h_{t}\in\Gamma_{i}$ be two distinct generators in the given code in $\mathrm{CAC^{e}}(p_{i},w)$, one has $j\cdot g_{t}\neq k\cdot h_{t}$. If $g_{t}=h_{t}$, it further implies that $(j-k)g_{t}=0$ (mod $p_{t}$), which is impossible because of $j,k$ beging distinct in $\{\pm 1,\pm 2,\cdots,\pm(w-1)\}$ and $p_{i}\geq 2w-1$. This completes the proof. ∎
@@
$K(p_{1}^{r_{1}}\cdots p_{t}^{r_{n}},w)=\frac{p_{1}^{r_{1}}\cdots p_{t}^{r_{t}}-1}{2(w-1)}.$ (39)
@@ V Multichannel CACs with Two Channels
Let $w$ be a positive integer, and let $p_{1}<p_{2}<\cdots<p_{n}$ be primes satisfying $p_{1}\geq w$. For each $1\leq i\leq n$, if there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $m_{i}$ codewords and the two conditions in (26) and (27) hold, then for any positive integers $r_{1},\ldots,r_{n}$, there exists a code $\mathcal{C}\in\text{MC-CAC}(2,L,w)$ with
@@
Since $\gcd(w-1,L^{\prime})=1$, we have $\mathbb{Z}\cong\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}$. Elements in $\mathcal{I}_{2}\times\mathbb{Z}_{L}$ can be represented by $(m,(z,x))$, where $m\in\mathcal{I}_{2}=\{1,2\}$ and $z\in\mathbb{Z}_{w-1}$, $x\in\mathbb{Z}_{L^{\prime}}$.
@@
$S_{a}^{1}$ $=\{(1,(1,-a))\}\cup\{(2,j(1,a)):\,j=0,1,2,\ldots,w-2\},\text{ and }$
$S_{a}^{2}$ $=\{(1,j(1,a)):\,j=0,1,2,\ldots,w-2\}\cup\{(2,(1,-a))\}.$
@@
We now show that the construction described in Theorem 11 is optimal.
@@
Let $w$ be positive integer, and let $p_{1}<\cdots<p_{n}$ be primes such that $p_{i}-1$ is divisible by 2w22w-2 for each $i$. For each $1\leq i\leq n$, if there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $(p_{i}-1)/(2w-2)$ codewords and the two conditions in (26) and (27) hold, then for any positive integers $r_{1},\ldots,r_{n}$,
@@
The assumption that $p_{1}-1$ is divisible by 2w22w-2 guarantees that $p_{1}\geq 2w-1\geq w$. By plugging $m_{i}=(p_{i}-1)/(2w-2)$ into Theorem 11, there exists a code in $\text{MC-CAC}(2,L,w)$ with $\prod_{i=1}^{n}p_{i}^{r_{i}}+\left(\left(\prod_{i=1}^{n}p_{i}^{r_{i}}\right)-1\right)/(w-1)$ codewords.
@@ VI Multichannel CACs with $M>2$ Channels
Let $M,L,L^{\prime}$, and $w$ be positive integers such that $w>M\geq 3$, $M$ divides $w$, $L=(2\frac{w}{M}-1)L^{\prime}$, and all prime factors of $L^{\prime}$ are larger than or equal to 2w12w-1. If there is a code in $\mathrm{CAC^{e}}(L^{\prime},w)$ with $(L^{\prime}-1)/(2w-2)$ codewords, then there exists a code $\mathcal{C}\in\text{MC-CAC}(M,L,w)$ with
$|\mathcal{C}|=\frac{M(M-1)L}{(2w-M)(w-1)}+\frac{M(L-1)}{2w-2}.$
@@
Let $\mathcal{C}^{\prime}$ be the code in $\mathrm{CAC}(L^{\prime},w)$ with $|\mathcal{C}^{\prime}|=(L^{\prime}-1)/2(w-1)$ codewords. For $S\in\mathcal{C}^{\prime}$ and $m\in\mathcal{I}_{M}$, define a $w$-subset in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ by
@@
Obviously, $|\mathcal{C}_{1}|=M(L^{\prime}-1)/(2w-2)$. It is easy to see that
@@
It is not difficult to see that
@@
Furthermore, by (64) and (66), it is obvious that $D_{S}\cap D_{S^{\prime}}=\emptyset$ whenever $S\in\mathcal{C}_{1}$ and $S^{\prime}\in\mathcal{C}_{2}$. It concludes that $\mathcal{C}=\mathcal{C}_{1}\uplus\mathcal{C}_{2}$ is a code in $\text{MC-CAC}(M,L,w)$. Finally, as $L^{\prime}=ML/(2w-M)$, the code size is
$|\mathcal{C}|$ $=|\mathcal{C}_{1}|+|\mathcal{C}_{2}|=\frac{M(L^{\prime}-1)}{2w-2}+L^{\prime}$
$=\frac{M+2w-2}{2w-2}\cdot\frac{ML}{2w-M}-\frac{M}{2w-2}$
$=\frac{M(M-1)L}{(w-1)(2w-M)}+\frac{M(L-1)}{2w-2},$
as desired. ∎

### v2

@@ IV-B Codeword Length $L=p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{n}^{r_{n}}$
This subsection is devoted to generalizing the following result, proposed in [20, Theorem 10].
@@
Let $w$ be a positive integer and let $p_{1}<p_{2}<\cdots<p_{n}$ be primes that satisfy $p_{1}\geq 2w-1$. If there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $m_{i}$ codewords for each $i$, then for any positive integers $r_{1},\ldots,r_{n}$, there exists a code $\mathcal{C}\in\mathrm{CAC^{e}}(L,w)$ with
@@
$g\in\mathcal{S}_{r_{i}}(\Gamma_{i}),x_{j}\in\mathbb{Z}_{p_{j}^{r_{j}}}\text{ for }i<j\leq n\}.$
@@
Suppose to the contrary that $d^{*}(S_{a})\cap d^{*}(S_{b})\neq\emptyset$, where $a=(0,\ldots,0,g,x_{i+1},\ldots,x_{n})$ and $b=(0,\ldots,0,h,y_{i+1},\ldots,y_{n})$ for some $g,h\in\mathcal{S}_{r_{i}}(\Gamma_{i})$. Assume $j(0,\ldots,0,g,x_{i+1},\ldots,x_{n})=k(0,\ldots,0,h,y_{i+1},\ldots,y_{n})$ for some $j,k\in\{\pm 1,\pm 2,\ldots,\pm(w_{i}-1)\}$. If $j=k$, then $g=h$ and $x_{t}=y_{t}$ for $i+1\leq t\leq n$, namely, $a=b$. So, we consider $j\neq k$ in what follows. Let us focus on the $i$-th component, namely, $jg=kh$. Since $p_{i}\geq 2w-1$, by Proposition 3, one has $jg\in\mathsf{L}_{t}$ for every $j\in\{\pm 1,\pm 2,\ldots,\pm(w-1)\}$ if $g\in\mathsf{L}_{t}$. Here we use $\mathsf{L}_{t}$ to denote the $t$-th layer in the $p_{i}$-ary representation of elements of $\mathbb{Z}_{p_{i}^{r_{i}}}$ for notatinoal convenience. It follows that $jg\neq kh$ for any $j,k\in\{\pm 1,\pm 2,\ldots,\pm(w-1)\}$ whenever $g$ and $h$ are in distinct layers. Therefore, we only need to consider the case when both $g$ and $h$ are in the same layer, say $\mathsf{L}_{t}$ for some $t$. Observe that $j,k\in\mathsf{L}_{0}$ due to $p_{i}\geq 2w-1$. By Proposition 3 again, $j\cdot g_{t}=k\cdot h_{t}$ (mod $p_{i}$). If $g_{t}\neq h_{t}$, by the assumption that $g_{t},h_{t}\in\Gamma_{i}$ be two distinct generators in the given code in $\mathrm{CAC^{e}}(p_{i},w)$, one has $j\cdot g_{t}\neq k\cdot h_{t}$. If $g_{t}=h_{t}$, it further implies that $(j-k)g_{t}=0$ (mod $p_{t}$), which is impossible because of $j,k$ being distinct in $\{\pm 1,\pm 2,\cdots,\pm(w-1)\}$ and $p_{i}\geq 2w-1$. This completes the proof. ∎
Example 2.
Let $w=4,n=2,p_{1}=17$, and $p_{2}=37$. One can check that $\Gamma_{1}=\{1,4\}$ forms a set of generators of a code in $\mathrm{CAC^{e}}(17,4)$ and $\Gamma_{2}=\{1,8,23,26,27,31\}$ forms a set of generators of a code in $\mathrm{CAC^{e}}(37,4)$. By Theorem 9, we have an equi-difference CAC of length $L=17^{r}\cdot 37^{s}$ and weight $4$ with $\frac{1}{8}(17^{r}-1)37^{s}+\frac{1}{6}(37^{s}-1)$ codewords, for any positive integers $r$ and $s$. When $r=s=1$, we have $\widehat{\Gamma}_{1}=\{(g,x)\in\mathbb{Z}_{17}\times\mathbb{Z}_{37}:\,g\in\Gamma_{1},x\in\mathbb{Z}_{37}\}$ and $\widehat{\Gamma}_{2}=\{(0,g)\in\mathbb{Z}_{17}\times\mathbb{Z}_{37}:\,g\in\Gamma_{2}\}$. The obtained code in $\mathrm{CAC^{e}}(629,4)$ has generators in $\{\theta^{-1}(a):\,a\in\widehat{\Gamma}_{1}\cup\widehat{\Gamma}_{2}\}$, where the bijection $\theta:\mathbb{Z}_{629}\to\mathbb{Z}_{17}\times\mathbb{Z}_{37}$ is given in (4). The generators obtained from $\{(1,x):\,x\in\mathbb{Z}_{37}\}$ are 1, 18, 35, 52, 69, 86, 103, 120, 137, 154, 171, 188, 205, 222, 239, 256, 273, 290, 307, 324, 341, 358, 375, 392, 409, 426, 443, 460, 477, 494, 511, 528, 545, 562, 579, 596, and 613. The generators from $\{(4,x):\,x\in\mathbb{Z}_{37}\}$ are given by adding $3$ modulo $629$ to those obtained from $\{(1,x):\,x\in\mathbb{Z}_{37}\}$. And, the generators from $\widehat{\Gamma}_{2}$ are 68, 119, 323, 408, 544, 578. The resulting CAC contains $80$ codewords.
We now show that the construction described in Theorem 9 is optimal in some cases.
@@
$K(p_{1}^{r_{1}}\cdots p_{n}^{r_{n}},w)=\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}-1}{2(w-1)}.$ (42)
@@
Note that equi-difference CACs of weight $w$ and length a prime $p$ with (p1)/(2w2)(p-1)/(2w-2) codewords are studied in [25, 26, 3, 33].
@@ V Multichannel CACs with Two Channels
By the help of constructions given in Theorems 6 and 9, we have the following construction for MC-CACs with $M=2$ channels.
@@
Let $w$ be a positive integer, and let $p_{1}<p_{2}<\cdots<p_{n}$ be primes satisfying $p_{1}\geq 2w-1$. For each $1\leq i\leq n$, if there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $m_{i}$ codewords and the two conditions in (27) and (28) hold, then for any positive integers $r_{1},\ldots,r_{n}$, there exists a code $\mathcal{C}\in\text{MC-CAC}(2,L,w)$ with
@@
Since $\gcd(w-1,L^{\prime})=1$, we have $\mathbb{Z}_{L}\cong\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}$. Elements in $\mathcal{I}_{2}\times\mathbb{Z}_{L}$ can be represented by $(m,(z,x))$, where $m\in\mathcal{I}_{2}=\{1,2\}$ and $z\in\mathbb{Z}_{w-1}$, $x\in\mathbb{Z}_{L^{\prime}}$.
@@
$S_{a}^{1}$ $=\{(1,-(1,a))\}\cup\{(2,j(1,a)):\,j=0,1,2,\ldots,w-2\},\text{ and }$
$S_{a}^{2}$ $=\{(1,j(1,a)):\,j=0,1,2,\ldots,w-2\}\cup\{(2,-(1,a))\}.$
@@
Example 3.
Let $w=4,n=2,p_{1}=7$, and $p_{2}=23$. We have $L=483$. It has been shown in Example 1 that $7$ and $23$ satisfy the two conditions in (27) and (28). It is easy to see that $\Gamma_{1}=\{1\}$ generates a code in $\mathrm{CAC^{e}}(7,4)$ and $\Gamma_{2}=\{1,5\}$ generates a code in $\mathrm{CAC^{e}}(23,4)$. By Theorem 11, we have a code in $\text{MC-CAC}(2,3\cdot 7^{r}\cdot 23^{s},4)$ with $7^{r}\cdot 23^{s}+\frac{1}{3}(7^{r}-1)23^{s}+\frac{2}{11}(23^{s}-1)$ codewords, for any positive integers $r$ and $s$.
Consider $r=s=1$. The first $7\cdot 23-1$ codewords (collected in $\mathcal{C}_{1}$) are those in the form
$\{(1,-(1,a)),(2,(0,0)),(2,(1,a)),(2,2(1,a))\}\quad\text{and}$
$\{(1,(0,0)),(1,(1,a)),(1,2(1,a)),(2,-(1,a))\}$
for all elements $a$ in $\widehat{Q}$, which is given in Example 1, and the ordered pairs (1,a),(1,a)(1,a),-(1,a) are considered in $\mathbb{Z}_{3}\times(\mathbb{Z}_{7}\times\mathbb{Z}_{23})$. Take $a=(1,21)\in\widehat{Q}_{1}$ as an example. As $a=113\in\mathbb{Z}_{161}$, $(1,a)=274\in\mathbb{Z}_{483}$ and $(1,a)=109\in\mathbb{Z}_{483}$ by the CRT correspondence, the two corresponding codewords are $\{(1,109),(2,0),(2,274),(2,65)\}$ and $\{(1,0),(1,274),(2,65),(2,109)\}$.
The $50$ codewords in $\mathcal{C}_{2}$ are those in the form
$\{(m,(0,0)),(m,(0,a)),(m,(0,2a)),(m,(0,3a))\}$
for $m=1,2$ and elements $a\in\widehat{\Gamma}=\{(1,x)\in\mathbb{Z}_{7}\times\mathbb{Z}_{23}:\,x\in\mathbb{Z}_{23}\}\cup\{(0,1),(0,4)\}$. Take $a=(1,21)$ as an example. As $(0,a)=435\in\mathbb{Z}_{483}$ by the CRT correspondence, the two corresponding codewords are $\{(1,0),(1,435),(1,387),(1,339)\}$ and $\{(2,0),(2,435),(2,387),(2,339)\}$. Finally, the last codeword (in $\mathcal{C}_{3}$) is $\{(1,0),(1,322),(1,161),(2,0)\}$.
@@
We now show that the construction described in Theorem 11 is optimal in some cases.
@@
Let $w$ be a positive integer, and let $p_{1}<\cdots<p_{n}$ be primes such that $p_{i}-1$ is divisible by 2w22w-2 for each $i$. For each $1\leq i\leq n$, if there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $(p_{i}-1)/(2w-2)$ codewords and the two conditions in (27) and (28) hold, then for any positive integers $r_{1},\ldots,r_{n}$,
@@
The assumption that $p_{1}-1$ is divisible by 2w22w-2 guaranties $p_{1}\geq 2w-1\geq w$. By plugging $m_{i}=(p_{i}-1)/(2w-2)$ into Theorem 11, there exists a code in $\text{MC-CAC}(2,L,w)$ with $\prod_{i=1}^{n}p_{i}^{r_{i}}+\left(\left(\prod_{i=1}^{n}p_{i}^{r_{i}}\right)-1\right)/(w-1)$ codewords.
@@
For $w=4$, the primes that satisfy the conditions given in Theorem 13 are 7, 607, 631, 751, 919, 1087, 1447, 2239, 2287, 2311, 2647, 2719, and so on.
@@ VI Multichannel CACs with $M>2$ Channels
For $M>2$ channels, we only focus on the scenario that weight $w$ is divisible by $M$. We first propose the following consruction.
@@
Let $M,L,L^{\prime}$, and $w$ be positive integers such that $w>M\geq 3$, $M$ divides $w$, $L=(2\frac{w}{M}-1)L^{\prime}$, and all prime factors of $L^{\prime}$ are larger than or equal to 2w12w-1. If there is a code in $\mathrm{CAC}(L^{\prime},w)$ with $h$ codewords, then there exists a code $\mathcal{C}\in\text{MC-CAC}(M,L,w)$ with $|\mathcal{C}|=Mh+L^{\prime}$ codewords.
@@
Let $\mathcal{C}^{\prime}$ be the code in $\mathrm{CAC}(L^{\prime},w)$ with $|\mathcal{C}^{\prime}|=h$ codewords. For $S\in\mathcal{C}^{\prime}$ and $m\in\mathcal{I}_{M}$, define a $w$-subset in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ by
@@
Obviously, $|\mathcal{C}_{1}|=Mh$. It is easy to see that
@@
Clearly, $|\mathcal{C}_{2}|=L^{\prime}$. It is not difficult to see that
@@
Furthermore, by (67) and (69), it is obvious that $D_{S}\cap D_{S^{\prime}}=\emptyset$ whenever $S\in\mathcal{C}_{1}$ and $S^{\prime}\in\mathcal{C}_{2}$. We conclude that $\mathcal{C}=\mathcal{C}_{1}\uplus\mathcal{C}_{2}$ is a code in $\text{MC-CAC}(M,L,w)$. ∎
Example 4.
Let $w=6,M=3$, and $L^{\prime}=37$. We have $L=3\cdot 37$. One can check that $\{0,1,2,3,4,5\}$ and $\{0,6,12,18,24,30\}$ form a code $\mathcal{C}^{\prime}\in\mathrm{CAC}(37,6)$. By Theorem 14, we have a code in $\text{MC-CAC}(3,111,6)$ with $2\cdot 3+37=43$ codewords. By the CRT correspondence $\theta:\mathbb{Z}_{111}\to\mathbb{Z}_{3}\times\mathbb{Z}_{37}$ given in (4), the first $6$ codewords are $\{(m,0),(m,75),(m,39),(m,3),(m,78)$, (m,42)}(m,42)\}, $\{(m,0),(m,6),(m,12),(m,18),(m,24),(m,30)\}$ for $m=1,2,3$. The remaining $37$ codewords are $\{(1,0),(1,a),(2,2a),(2,3a),(3,4a),(3,5a)\}$, where $a\in\Omega=\{(1,g)\in\mathbb{Z}_{3}\times\mathbb{Z}_{37}:\,g\in\mathbb{Z}_{37}\}$. The set $\Omega$ consists exactly of the multiples of 3 in $\mathbb{Z}_{111}$.
Note that the base CACs for the construction in Theorem 14 are not necessarily equi-difference nor tight.
Corollary 2.
If the based CAC described in Theorem 14 has $(L^{\prime}-1)/(2w-2)$ codewords, then there exists a code $\mathcal{C}\in\text{MC-CAC}(M,L,w)$ with
$|\mathcal{C}|=\frac{M(M-1)L}{(2w-M)(w-1)}+\frac{M(L-1)}{2w-2}.$ (70)
Proof.
The result is obtained by plugging $L^{\prime}=ML/(2w-M)$ and $h=(L^{\prime}-1)/(2w-2)$ into $|\mathcal{C}|=Mh+L^{\prime}$. ∎
The code size in (70) coincides with the upper bound given in Theorem 4, namely (16). However, the code constructed in Corollary 2 is not optimal since its length contains a factor, say $2\frac{w}{M}-1$, that is smaller than 2w12w-1. Therefore, Theorem 4 is not applicable in this case. In the following, we derive another upper bound for CACs of this length.